VIBRATIONAL SPECTROCOPY OF THE OCCN RADICAL THROUGH TIME RESOLVED FOURIER TRANSFORM INFRARED EMISSION SPECTROSCOPY

\underline{W. MCNAVAGE} and H. L. DAI, Department of Chemistry, The University of Pennsylvania, PA 19104.

The CN stretch of the OCCN radical was observed for the first time at 2093 cm^{−1}. The method of detection was by nanosecond Fourier-transform infrared emission spectroscopy. The radical was produced through UV photodissociation(λ=193 nm) of carbonyl cyanide, CO(CN)_{2} and pivaloyl cyanide, CO(CN)(CH_{3})_{3} which leaves the radical with sufficient internal excitation. Infrared and near-infrared emission from all vibrationally excited species was detected. Pressure dependence, rotational band contour analysis and ab initio calculations all aided in assignment.